• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name N Toro

    Log entry time 07:59:56 on July 9,2010

    Entry number 331024

    keyword=Hall A Tools and magnet stripchart at beginning of run 1934 -- Cycling left quads to 1.16 (right quads fixed since last cycling)

    Now cycling HRS-L Q2 and Q3 (see stripchart).
    HRS-R has not been changed since last cycled to 1.16 (also shown in stripchart).

    Will now take sieve-out data with both spectrometers at 1.16.

    (The reason for this complicated procedure was that we took data at -4% in the left arm only while waiting for the right dipole to settle down.)


    FIGURE 1

    FIGURE 2